Consider two 100microfarads capacitors.one charged to 10 V, disconnected from power supply and then connected across the other.calculate the energy stored by the combination


1
Expert's answer
2021-03-29T08:58:00-0400
"q=CU\\C'=2C\\
E=0.5\frac{q^2}{2C'}\\
E=0.5\frac{C^2U^2}{4C}\\
E=\frac{CU^2}{8}\\
E=\frac{(10^{-4})(10)^2}{8}=1.25\ mJ"